OpenStudy (anonymous):

On the first day of a road trip, you travel 412 miles on one tank of gas. The distance you travel varies directly with the amount of gas you use. If your gas tank can hold 16 gallons of gas, approximately how many total gallons will you use if the total distance you are traveling is 1150 miles? Select one: a. 30 b. 45 c. 55 d. 32

OpenStudy (ahsome):

I have this feeling that the total distance travelling is more than 1150 miles? Please conform @Yolo4mecuite

OpenStudy (anonymous):

Yes

OpenStudy (ahsome):

The I am not sure. It seems that the possible answers are too high, compared to the fact that one single gallon can get 412 miles

OpenStudy (anonymous):

OK.

OpenStudy (campbell_st):

it is simply a case of calculating the fuel consumption in miles per gallon Fuel Consumption = 412/16 miles per gallon Fuel for your journey = distance/fuel consumption. Hope it helps

OpenStudy (anonymous):

I got 44.66 so rounded 45?

OpenStudy (campbell_st):

an alternate method is Tanks of Fuel = 1150/412 then fuel = tanks of Fuel * 16